king about opening a soccer camp in southern California. To start the camp, Coolplay would need to purchase land and build Coolplay Corp. is thi four soccer fields and a sleeping and dining facility to house 150 soccer players. Each year, the camp would be run for 8 sessions of 1 week each. The company would hire college soccer players as coaches. The camp attendees would be male and female soccer players ages 12-18. Property values in southern Calfornia have enjoyed a steady increase in value. It is expected that after using the facility for 20 years, Coolplay can sell the property for more than it was originally purchased for. The following amounts have been estimated. $324,000 Cost of land $648,000 Cost to build soccer fields, dorm and dining facility Annual cash inflows assuming 150 players and 8 weeks $993,600 $907.200 Annual cash outflows 20 years Estimated useful life $1,620,000 Salvage value
Assume that during the first 5 years, the annual net cash flows each year were only $43,200. At the end of the fifth year, the company ls running low on cash, so management decides to sell the property for $1,438,560. What was the actual internal rate of return on the project? (Round answer to 0 decimal places, 4.13%. For calculation purposes, use 5 decimal places as displayed in the factor table provided.) Actual internal rate of return
Answer #1
NPV AT 8%
CASH FLOWS ANNUITY AT 8% PRESENT VALUE OF CASH FLOWS
YEAR 0 [324000+648000]972000 1 -972000
YEAR 1- 5 43200

3.99263

=PV[1/1.08]5

=(0.9259+0.85733+0.7938+0.7350+0.6806)

172481

[43200*3.99263]

YEAR 5 1438560 0.6805 [1/1.08]5 978940
NPV

179421

[-972000+172481+978940]

IRR IS THE RATE OF RETURN AT WHICH NPV BECOMES ZERO.

WE WILL FIND OUT NPV AT 12%

NPV = -972000 + 43200 FV[ 1/1.12]5 + 1438560 [1/1.12]5

          = -972000+ 43200[3.6046] +1438560 [0.5674]

=-972000+155718+816239

NPV = - 43

IRR = R1 + [(NPV 1 * (R2 - R1)) / (NPV1 - NPV 2)]

R1 = RANDOMLY SELECTED RATES OF RETURNS 8%

R2 = RANDOMLY SELECTED RATE OF RETURN 12%

NPV1 = 179421

NPV 2 = -43

IRR = 0.08 + [(179421 * (0.12-0.08) / (179421 - (-43)]

=0.08 + [7176.84 / 179464]

=0.08+0.0399

=11.99%

ANSWER = 12% ROUNDED OFF.